The joint probability density function for continuous random variables X and Y is given below.

f (x) =   x + y, 0 < x < 1, 0 < y < 1 if;

                                                                         0,         degilse.

(a) Show that this is a joint density function.

(b) Find the marginal density of X .

(c) Find the marginal density of Y .

(d) Given Y = y find the conditional density of X . (e) P ( 1/2 < X < 1|Y = 1/2 ) =?                               

(f ) Are these random variables statistically independent?
Answer #1

Solcetion: f(x,y) = +y, Oceal, ocycl a) f(x,y) dy da al X xty) dy dx - y] + you, ax 5 (13) ó ) dx = [***.+ 23:] 2+1 dx = =of (a) marginal J dx s fror, 4) fylyla (octy ) dx 3. 9=1, y + b ocyel f(x,y) 20+y (d) Ocael f (xcly) = fy(y) y+12 ocyel e) P